Reading List Gregor von Laszewski PhD Quallification Exam

1 Research Statement

• Name: Gregor von Laszewski

• Project Leader: Miloje Makivic, Kim Mills

• Thesis Supervisor: Geoffrey Fox

• Academic Supervisor: Geoffrey Fox

• Research Statement My research area is “Heterogeneous Supercomputer Environments” which combines the use of SIMD and MIMD algorithms to solve real application problems. The project is in tight connection with the NASA project about the parallelization of an optimal interpolation algorithm. The computational demands of this program are to be expected so high that a Heterogeneous computing environment is useful for solving the problem.

• Education Plan:

Qualification exam: first quarter 94 Expected Graduation: summer quarter 95

• Exception Statement: N/A

• Previous NPAC agreements:

The role of the academic advisor has been fulfilled by Sanja Ranka. The Roll of the Project Leader has been changed from A. Mohamed to Thomasz Haupt to Kim Mills.

• Exam Commitee Members:

Geoffry C. Fox, Sanjay Ranka, Alok Choudhary

2 Readinglist

General Parallel [?] [?] [?] [?] [?] Parallel Climate Modeling and Load Balance Issues [?] [?] [?] [?] [?]

[?] Static Load Balance Sequential Algorithms [?] [?] [?]

Parallel Algorithms [?] [?] [?] Simulated Annealing [?] [?] Genetic Algorithm [?] [?] [?]

Task Scheduling [?] [?] [?] [?] [?]